OpenStudy (anonymous):

Eliminate the “z” variable in the last two equations by finding opposite coefficients and adding. What is the resulting equation? x + 3y + 2z = –16 2x – y + 2z = –15 2x – 2y – z = 1

OpenStudy (anonymous):

4x + y = –14 6x – 5y = –13 6x + 5y = 13 4x – 3y = –14

OpenStudy (anonymous):

okay i got 6x – 5y = –13